How many kg of ore containing 5% gold must be melted down with 8kg of 2% gold to produce a mixture which is 4% gold?
Answer by mananth(16946)

-------------- percent ---------------- Amount
Ore I 2 ---------------- 8 kg
Ore II 5 ---------------- x kg
Mixture 4 ---------------- 8 + x kg

2 * 8 + 5 x = 4 ( 8 + x )
16 + 5 x = 32 + 4 x
5 x -4 x = -16 + 32
1 x = 16
/ 1
x = 16 kg Ore II

m.ananth@hotmail.ca